OpenStudy (teller):

The Cooper family and the Nelson family each used their sprinklers last summer. The Cooper family's sprinkler was used for 35 hours. The Nelson family's sprinkler was used for 40 hours. There was a combined total output of 1325L of water. What was the water output rate for each sprinkler if the sum of the two rates was 35L per hour?

OpenStudy (anonymous):

x+y=35 35x+40y=1325 Know how to solve this system of equation?

OpenStudy (teller):

No I don't

OpenStudy (anonymous):

Solve for y by canceling x You do this by multiplying x+y=35 by (-35)

OpenStudy (teller):

Huh???

OpenStudy (teller):

I'm a little confused sorry...

OpenStudy (anonymous):

 Know what to do next? You should have known how to solve this equation. Did you do your other assignment involving this? This requires two steps to solve so it is harder than just the second ste?

OpenStudy (teller):

We don't have lessons ... were in a class on a website but it's through another program you just complete the work it gives you like three to four questions but if you get one wrong it sends you back to the start...

OpenStudy (teller):

I just need help on this .

OpenStudy (anonymous):

You should have a done a problem similar to this equation.

OpenStudy (teller):

Nope

OpenStudy (teller):

his is the first question ....

OpenStudy (teller):

This

OpenStudy (anonymous):

The x and y on both equation are the same. x is Cooper water output rate y is Nelson water output rate.

OpenStudy (anonymous):

Now see if you can solve it

OpenStudy (teller):

Yes but how do you solve it....

OpenStudy (teller):

Your basically not telling me how to solve it like I don't know how to set it up or anything like that..

OpenStudy (anonymous):

Read what I wrote above. You solve for x and y. Got to reread the whole thing over...

OpenStudy (anonymous):

Read the question again too, so you can better understand it and what I am saying with it.